ganeshie8 (ganeshie8):

Challenge problem : Show that \(F(a) = \frac{2^{2^{a+1}}-1}{2^{2^a}+1}\) is an integer for all \(a \in \mathbb{N} \)

ganeshie8 (ganeshie8):

\[\large F(a) = \frac{2^{2^{a+1}}-1}{2^{2^a}+1} \]

OpenStudy (bohotness):

i think i can answer it

ganeshie8 (ganeshie8):

@bohotness im sure you can... please go ahead :)

OpenStudy (bohotness):

i have it on my phone and i try to frigre the rest out

OpenStudy (bohotness):

but i guess you need to find out what a is

OpenStudy (bohotness):



OpenStudy (freckles):

\[2^{2^1}+1=2^2+1=4+1=5 \\ 2^{2^{1+1}}-1=2^{2^{2}}-1=16-1=15 \\ \text{ and looks good for } a=1 \text{ since } 5|15\]

OpenStudy (bohotness):

2^2(a+1)-1

ganeshie8 (ganeshie8):

suppose \(\large 2^{2^a}+1~ |~2^{2^{a+1}}-1 \) we need to show \(\large 2^{2^{a+1}}+1~ |~2^{2^{a+2}}-1 \)

OpenStudy (bohotness):

k

OpenStudy (bohotness):

2*l+220*i+221

OpenStudy (zarkon):

difference of squares

ganeshie8 (ganeshie8):

that will do haha! xD

ganeshie8 (ganeshie8):

\[\large F(a) = \frac{2^{2^{a+1}}-1}{2^{2^a}+1} = \frac{2^{2\cdot {2^{a}}}-1}{2^{2^a}+1} = \frac{\left(2^{ {2^{a}}}\right)^2-1}{2^{2^a}+1} \] not realy a challenging one :P:

OpenStudy (bohotness):

wow

ganeshie8 (ganeshie8):

A straightforward induction also requires using something like this i guess

OpenStudy (bohotness):

oo k

ganeshie8 (ganeshie8):

Induction step for freckles could be like this :\[\large 2^{2^{a+2}}-1 = \left(2^{2^{a+1}}\right)^2-1 = (2^{2^{a+1}}+1)(2^{2^{a+1}}-1)\] I don't see a way out if the difference of squares identity doesn't strike...

OpenStudy (kainui):

Fun =D How did you find this, or was the whole thing just supposed to be a tricky way of writing a difference of two squares? xD

ganeshie8 (ganeshie8):

Hahaha it just messes with difference of squares ;) saw this neat result in proving \(2^{2^a}+1\) when prime has no primitive roots : \[2^{2^{a+1}}-1 = \left(2^{2^{a}}\right)^2-1 = (2^{2^{a}}+1)(2^{2^{a}}-1)\] That means \(2^{2^{a+1}}-1\equiv 0 \pmod{2^{2^{a}}+1}\) However the number of coprime integers less than \(2^{2^{a}}+1\) would be \(2^{2^a}\) which is always greater than \(2^{a+1}\) proving no primitive roots

OpenStudy (anonymous):

:D